OpenStudy (anonymous):

**Please help** A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 4 experience points. On level 3, a player must earn 8 experience points. On level 4, a player must earn 13 experience points. Which recursive equation represents the pattern? an = an – 1 + n an = an – 1 + (n + 1) an = an – 1 + 2n an = an – 1 + (2n + 1)

OpenStudy (anonymous):

looks like each level you have to earn 2 + the number the last level was in order to advance

OpenStudy (anonymous):

Yes,

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

n is the number of the level

OpenStudy (anonymous):

'an' is the number of experience points you need to pass the nth level

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

To be honest, I tried understanding this a month ago and I don't need to understand it anymore. Can you just tell me the answer?

OpenStudy (anonymous):

look at my first comment now

OpenStudy (anonymous):

if you don't need to understand it, why are doing this problem?

OpenStudy (anonymous):

Re-submitting a practice test I did months ago.

OpenStudy (anonymous):

I already finished this lesson and what not. I just need the answer

OpenStudy (anonymous):

then look at my first comment, it basically tells you the answer

OpenStudy (anonymous):

an = an – 1 + 2n?

OpenStudy (anonymous):

That was wrong...

OpenStudy (anonymous):

if you're on level 2, you need 4 experience points so plug in your values and see if that's true 4 = 4-1 + 2*2 4 = 3+4 4 doesn't equal 7, so no, that's not the correct answer